Find the derivatives of the following functions.
1. f(t) = t^11
2. g(t) = 6√t
3. g(w) = 5/6w^12
4. Let f(t) = t^3 − 27t + 5
a. Find the value of t for which the slope of the curve y = f(t) is 0.
b. Find the value of t for which the slope of the curve y = f(t) is 2t.
5. Find f^0(x), f^00(x), f^000(x) for the following function.
f(x) = 3x^12 + 4x^3


Answer by Alan3354(69443)

Find the derivatives of the following functions.
1. f(t) = t^11
===========
2. g(t) = 6sqrt(t)
g(t) = 6*t^(1/2)
--------------------
3. g(w) = 5/6w^12
===================
4. Let f(t) = t^3 − 27t + 5
a. Find the value of t for which the slope of the curve y = f(t) is 0.
b. Find the value of t for which the slope of the curve y = f(t) is 2t.
5. Find f^0(x), f^00(x), f^000(x) for the following function.
f(x) = 3x^12 + 4x^3